OpenStudy (anonymous):

solve the equation cos^2x-sin^x=-cosx

OpenStudy (anonymous):

dude, http://www.wolframalpha.com/input/?i=cos%5E2x-sin%5Ex%3D-cosx

OpenStudy (anonymous):

$$\cos^2 x-\sin^2x=-\cos x$$Recognize that \(\cos^2x-\sin^2x=\cos2x\) and thus we have$$\cos2x=-\cos x$$So now you can interpret \(x\) as the angle s.t. if you double it you get get reflected into the opposite quadrant: